OpenStudy (anonymous):

Expand the binomial. (5 − y)4

OpenStudy (anonymous):

(5 − y)^4

jhonyy9 (jhonyy9):

(5-y)^2(5-y)^2 =(25-10y+y^2)(25-10y+y^2)

jhonyy9 (jhonyy9):

- do you need to continue ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

and can you make it ?

OpenStudy (anonymous):

sorry im back no i dont know how

OpenStudy (anonymous):

This the binomial Theorem: \[(x+y)^n=\sum_{k=0}^{n} \left(\begin{matrix}n \\ k\end{matrix}\right) x^{n-k}y^k\]